A Complete Bibliography of Publications in History of Geo- and Space Sciences

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

15 April 2025 Version 1.15

Title word cross-reference

27 [HDPG^{+}13].

-day [HDPG^{+}13].

1 [Hul11, Lin23a]. 100th [FS12]. 1767 [Apl19b]. 1770 [CAOV18, Sch10]. 18 [Sch10]. 18th [SS11]. 1910 [Apl19a]. 1915 [dVKR^{+}23]. 1950-1966 [ZW21]. 1960s [Dam18]. 1980 [Car14]. 19th [Gou22, NH12].

2 [Jør11, Lin23b, Oks11, Sta11b]. 2011 [Jør11, Sta11b]. 20th [GLMMLGGA25, LGGAGLdA19]. 21st [JIZB^{+}19].

3 [Hol12, Lin24].

4 [Hae16].

6 [MFN23].

80th [Har11].

academia [CH20]. Academy [ZW21]. account [Vaq17]. Actions [ZC18]. Adriatic [Rai20]. advanced [SL14]. aeromagnetic [RSD^{+}13]. aeronomy [SG11]. Africa [Kot18, Mar16, Med13]. After [ML16]. again [ML16]. Agassiz [Ate22]. age [Ate22]. airship [RSD^{+}13]. Alec [LGGAGLdA19]. Alexander [MKSG10]. Alibag [GDN15]. Almstedt [Gla20]. America [Aal15]. Amundsen [Dee11]. ancient [Mar12, Sch24, Shc18].

Andean [Riv11, Car10]. Andøya [Thr18]. anniversary [FS12, LGS^{+}20]. Anthropic [Cat11]. Applications [BMvRS12, Wil15]. approach [San17]. approaches [Ant22]. Arc [Puz19]. Arctic [RSD^{+}13]. Arecibo [Mat13]. Arge [She17]. Arrhenius [Kra13]. Arthur [GALGGL18]. article [BJB11]. artist [BN15]. ASFC [WKC18]. aspects [DPR12]. assessment [MKSG10]. Association [AFH^{+}19, DÁ19, Cas19, MFN23]. atmosphere [CW14]. Atmospheric [Apl18, HR22, HS23, HR24, AH13, BSB^{+}20, FS17, Har20, MV19, Apl20]. attempt [Ert11]. attraction [Sma18]. Augustus [BN15, Nat16]. aurora [Ege24, Sch10, Sta11c, CAOV18, Sil12]. Auroral [EB12, Bur16, EB19, Nag13, FVZF14, Jør11, Kra13, MS12, NT24, SS11, SL14, Sta11a, Sta11b]. auroras [FGFL12]. Australian [WKC18]. Austria [Len21].

B [BJB11]. Ball [Vaq17, DC18, Keu21]. based [MKSG10, Row23, ZW21]. Bauer [Har11]. Became [Mil09, Sch13]. before [dVKR^{+}23]. beginning [IZJ19]. beginnings [Cas19]. Behind [Har20, Cha18, Ekm16]. belief [Edw16, Sud14]. Bernhard [HDPG^{+}13]. between [Ert11, ZC18, Har11]. bicentenary [GMdSM21]. biography [Bur12]. Birkeland [EB10]. birth [JK16]. birthday [Har11]. Bjerrum [GLMMLGGA25]. body [DÁ19]. Book [Apl19a, Apl19b, Bre16, Car14, Cha18, Lan24, Liu24, Nag13, Nat16, Ric12, Riv11, Sch13, Sil12, Tay10, Wil15]. brief [Ate22, Cat11, HWD^{+}18, Keu21, MH13, RR19, Wit20]. Burt [Apl19b, Apl19b].

calendar [Gaj19]. Carl [EB12, GT14, Wit20, Nag13]. Carnegie [Har20]. cartographer [BN15]. Casagrande [GALGGL18]. case [Gaj19, San17]. Celsius [Cha18, Ekm16]. cementation [HGTJ21]. centenary [IR17]. Center [Dam18, HWD^{+}18]. centers [Wil18]. Central [HGTJ21]. Centre [WKC18]. Century [Ric12, Bar15, GALGGL18, Gou22, GLMMLGGA25, JIZB^{+}19, LGGAGL19, LGGAGLdA19, MV19, NH12, RTR11, SS11]. Challenger [Gou22]. Chalmers [Apl18]. Change [Car10, DÁ19, Riv11]. changes [Ekm16, Cha18]. chapter [Kra15]. charts [MKSG10]. Chatanika [MH13]. Chérifien [Med13]. China [HWD^{+}18, ZW21]. Chinese [ZW21]. Christian [SS11, NH12]. circuit [Har20]. Climate [Apl19b, Car10, Riv11, Bar15]. clocks [Agn20]. clouds [DPR12]. co [Gla20]. co-founder [Gla20]. coast [Mar16]. coastlines [Shc18]. Colaba [GDN15]. Colonial [Car14, Ert10]. commencement [SN21]. Comment [Jør11, Edw16, Sta11b]. Commission [LASJ22]. common [JIZ19b]. community [ZC18]. compared [Gou22]. comparison [Har20]. completely [Ert11]. concepts [CKC^{+}22]. configuration [Shc20]. confluence [And21]. Conjugate [Ege24]. contents [Pis14]. context [Sch23]. continuity [And21]. continuous [SBLW14]. Contribution [GALGGL18, GLMMLGGA25, LGGAGL19, LGGAGLdA19, Med13, Hae16, HVR23, LGS^{+}20, Val17]. contributions [Apl18, Dee11, Nis10]. Cook [WR18]. cooperation [MV19]. coordinates [Mar11, Mar12]. copper [HGTJ21]. cosmic [FS12]. cosmical [Kra13]. Cosmogenic [BMvRS12, Wil15]. COST [ZC18]. creation [Med13]. Cryospheric [AFH^{+}19, BJB11]. Cummings [Liu24]. curious [Gaj19]. current [Bur12]. Currents [Mil09, Sch13]. curve [Har20]. cycle [CKC^{+}22]. cycles [HDPG^{+}13]. Cyclicity [CKC^{+}22]. cycling [HCB^{+}10]. Czech [HS21].

Danish [Jør11, Sta11b, NH12, Sta11a, Sta11c]. Danmarks [Sta11c]. Darmstadt [Dam18]. data [dVKR^{+}23]. David [Liu24]. day [HDPG^{+}13]. days [dVKR^{+}23]. death [IR17]. Debates [Liu24]. decades [Wan22]. declination [ML16, TPCNA24]. declining [Kra15]. deformation [Car21]. Degree [Cha18]. Degrees [Ekm16]. density [Sma18]. designer [BN15]. Deutsche [Sch23]. developer [NH12]. development [BJB11, EZ13, FS12, IZJ19, Med13, Puz19, Sch10, Val17, Wan22, WKC18, ZW21]. different [Ant22, JIZ19b]. Dioscuri [FS17]. directorship [Bur16]. discharges [Ert11]. discovered [SN21]. discoveries [Mad23]. discovery [FS12, Her10, Ege24]. discrete [FVZF14]. distribution [FVZF14]. disturbances [EB10]. Dobrowolski [BJB11]. Dr. [Har11, SG11]. draughtsman [BN15]. drift [dVKR^{+}23]. Düll [HDPG^{+}13]. Durham [Apl18]. during [GALGGL18, GLMMLGGA25, LGGAGL19, LGGAGLdA19, Wan22, WR18, dVKR^{+}23]. dynamics [FVZF14].

earliest [DPR12]. Early [NT24, Nis10, SN21, Car21, DC18, GALGGL18, GLMMLGGA25, Hae16, Hol12, Hul11, IZJ19, LGGAGL19, LGGAGLdA19, Nev14, Oks11, Ric12, SS11, SL14, Spe21, Win23, WFBM19, ZW21]. Earth [Cha18, Edw16, Ozc20, CKC^{+}22, Edw16, Dee11, Ekm16, Kra15, MKSG10, Sma18, Sud14]. earthquake [SBLW14]. East [Car14, Ert10]. Ebstorf [Pis14]. Editorial [AAE^{+}10]. Eduard [Spe21]. education [JSBWL16]. Edwin [Aal12]. Egeson [HCB^{+}10]. Eighteenth [RTR11, Ric12]. Eighteenth-Century [Ric12, RTR11]. EISCAT [BGK^{+}13, Hae16, Hol12, Hul11, MFN23, Oks11, Wan22]. electric [Har20]. electrical [Apl20]. electricity [AH13, Apl18, BSB^{+}20, FS17, HR22, HS23, HR24].

Elster [FS17]. emissions [EB19]. Empire [OO14]. encyclopedias [Sch24]. Endeavour [WR18]. Endurance [dVKR^{+}23]. Envelope [Mil09, Sch13]. Environmental [Fro10]. Environments [BMvRS12, Wil15]. Erich [CW14]. Eskdalemuir [HR24]. Essay [CKC^{+}22]. established [CH20]. establishment [IZJ19]. estimate [Sma18]. Europe [HGTJ21]. European [ZC18]. Euxinus [Shc20]. eventful [Ric12]. Evolution [Apl19a, TPCNA24]. evolves [JIZ19a]. Expanding [Kra15]. expansion [Edw16, Sud14]. expeditions [Ege24]. experienced [Jac22]. explanations [Sch24]. Exploration [Lan24]. Explorer [Bre16, GMdSM21, Hai83]. Extreme [Sch15].

fabrics [Car21]. facility [RS22]. facts [SN21]. Father [PA12]. female [Jac22]. field [Ant22, FGFL12, MKSG10]. fields [Dee11]. Finland [Bös21, NT24, Oks11]. Finnish [Nev14]. first [BJB11, LASJ22, PA12, RSD^{+}13, Row23, Wan22]. flight [RSD^{+}13]. Fluid [Mil09, Sch13]. Forecast [WKC18]. formation [And21]. forms [FVZF14]. Foundation [Sch23, Win23]. founder [Gla20, IR17]. founding [EZ13]. fragile [Jac22]. Francisco [SBLW14]. Franz [Mei17]. French [BGK^{+}13]. Friedrich [GT14, Gla20, Wit20, IR17]. frontier [SWGM^{+}19]. Fürstenfeldbruck [Sof15]. future [AFH^{+}19, KSDJ19, MP19b].

G [Lan24]. Garmisch [HS23].

Garmisch-Partenkirchen [HS23]. gauge [Row23]. gauge-based [Row23]. Gauss [GT14, Wit20]. Gazelle [Gou22]. Geitel [FS17]. General [GT14]. Geo [Jør11, Sta11b]. Geodesy [JIZ19b, Edw16, IR17, DÁ19]. Geodetic [Puz19, HS21]. geographic [Mar11, Mar12]. Geographical [JSBWL16, San17].

Geographike [Mar11, Mar12]. Geography [Shc20, Mar16, Shc18]. geologic [Aal12, Aal15]. geological [CKC^{+}22]. geologist [GMdSM21, Mul24]. geologists [Ric12]. Geology [RTR11, Ric12]. Geomagnetic [Sof15, EB10, HVR23, Lin23a, Lin23b, Lin24, SWS10, TPCNA24]. geomagnetism [Kot18, SG11]. Geophysical [BSB^{+}20, Bös21, Gla20, HR22, HR24, LGS^{+}20, NT24, ACFJ13, Mat13, Len21]. Geophysics [JIZ19b, JK16, Kra15, OO14, ZW21, ZW21]. Geophysikalische [Sch23]. Georg [GMdSM21, SWS10]. George [HCB^{+}10]. Geoscience [IZ16, NH12]. geoscientific [Med13]. geotechnics [GALGGL18, GLMMLGGA25, LGGAGL19, LGGAGLdA19]. German [GT14, Gla20, GMdSM21, Hae16, Mei17]. Gesellschaft [Sch23]. gifted [Jac22]. Gjøa [Ege24]. Glaciers [Car10, Riv11]. Global [LASJ22, DÁ19, Har20, SL19]. globe [JIZ19b]. Göttingen [SBLW14]. Graf [RSD^{+}13]. gravimetric [Pet16]. gravimetry [Edw16]. gravity [Agn20, Ant22, Kra15]. Great [CAOV18]. Greece [Sch24]. group [LASJ22]. Günther [FS12].

Haerendel [Lan24]. Haldde [Bre21]. half [Bar15]. Hans [FS17]. Happiness [Car14, Ert10]. Harald [Sta11c, Sil12]. Harang [Bur16]. Hartung [GMdSM21]. heating [RS22]. Hector [Bre16, Hai83]. Helena [CWR17]. Heliogeophysical [KSDJ19]. Helliwell [CI12]. Helmert [IR17]. Hermann [Aal15]. Hermanus [Kot18]. Heuson [SS11]. high [Car21]. Hinton [Aal12]. Hist [Jør11, Sta11b]. Historian [SG11]. historic [EB19]. Historical [HVR23, TPCNA24, CKC^{+}22, DPR12, FVZF14, Har20, Kot18, Pet16, SN21].

History [EZ13, FN22, Hae16, Hol12, Hul11, Lin23a, Lin23b, Lin24, MFN23, Oks11, Ric07, RS22, Sof15, Tay10, Wan22, Wei16, Ate22, Cat11, HWD^{+}18, Her10, Jør11, Keu21, Kra15, Len21, Mat13, MH13, MP19b, Nev14, OO14, PB11, RR19, SN21, Sta11a, Sta11b, Thr18, WFBM19, ACFJ13, Wil18]. HMS [Gou22]. Hooke [Ric12]. human [HDPG^{+}13]. Humboldt [MKSG10]. Hungary [BSB^{+}20]. Hutton [Sma18]. hydrogen [EB19]. hydrological [Val17]. hydrology [RR19]. Hyphegesis [Mar11, Mar12]. hypothesis [Edw16, Sud14].

IACS [AFH^{+}19]. IAGA [MP19a]. IAHS [RR19]. IAMAS [MV19]. IAPSO [SWGM^{+}19]. IASPEI [SL19]. IAVCEI [Cas19]. ice [Ate22]. Iceland [Kri12]. ideal [DÁ19]. illustrator [GMdSM21]. importance [LASJ22]. improve [Sma18]. Incoherent [MP19b, BGK^{+}13, WFBM19]. Indications [Edw16]. Indies [Car14, Ert10]. industry [CH20]. influence [CKC^{+}22, GDN15]. Inge [Jac22]. inherently [Jac22]. initiation [Puz19]. Inseparability [Her10]. Institut [Med13, Sta11c]. Institute [CH20, EZ13, Nev14, Sta11c, Woo20, ZW21]. institutes [ACFJ13]. instrumentation [FS12]. instruments [HS21]. Intellectually [Jac22]. International [JIZ19b, Cas19, IZ16, LASJ22, MV19, Sch23, AFH^{+}19, DÁ19, LGS^{+}20]. Introduction [Apl20, ACFJ13, PB11]. Introductory [AAE^{+}10]. Investigating [Cha18, Ekm16]. investigation [SN21]. Investigations [FVZF14, EB10]. investigators [FS17]. involvement [BGK^{+}13]. ionisation [CW14]. ionosphere [RS22, EZ13]. ionospheric [And21, PB11, SL14, Wei16, ZC18]. Irkutsk [MP19b]. irradiance [Sch15]. irregular [DÁ19]. Irrigation [Car14, Ert10]. Islamic [Ozc20]. issue [Apl20, JIZ19b]. Istanbul [Ozc20]. István [BSB^{+}20]. Italian [PA12]. IUGG [IZJ19, JIZB^{+}19, JIZ19a].

J [Har11, Liu24]. J. [Apl18]. Jagiellonian [JSBWL16]. James [Aal12, Hai83, WR18, Bre16]. January [Sch10, CAOV18]. Japan [MFN23]. Japanese [Nis10]. Javanese [Ert11]. Jicamarca [WFBM19]. Johann [SS11]. John [Aal12]. Johnson [CWR17]. journey [Ozc20]. journeys [HGTJ21]. Julius [FS17]. Juliusruh [Wei16].

Kakioka [FN22]. Karl [Gla20, RS18]. Karsten [Aal15]. Kelvin [AH13]. Kew [Apl19a]. knowledge [Dee11, GALGGL18, GLMMLGGA25, LGGAGL19, LGGAGLdA19]. known [DPR12, SS11]. Koch [Nat16, BN15]. Kolupaila [Val17]. Kossmat [Mei17]. Kristian [EB10].

Lanzerotti [Liu24]. large [Mad23]. large-scale [Mad23]. last [dVKR^{+}23]. latitude [San17]. Laurits [GLMMLGGA25]. Lavoisier [Ric12]. Leader [Bre16]. legacy [Apl18, Bre21, Kri12]. legendary [Mul24]. Lehmann [Jac22]. Leibniz [Ric12]. length [Shc18]. Lerwick [HR22]. level [Rai20, Row23]. Life [Lan24, BN15]. Lightning [Sch24, DC18, Keu21, Vaq17]. Lights [Bur16]. little [SS11]. little-known [SS11]. Liverpool [CH20, Woo20]. Locales [Ert10, Car14]. location [Mar16]. Lord [AH13]. Louis [Liu24]. luminosity [FVZF14].

machines [Woo20]. Maclure [Aal12]. Madrid [TPCNA24]. Magnetic [FN22, Kot18, Dee11, FGFL12, GDN15, MKSG10, MP19a, PA12]. Magnetism [GT14]. Maisach [Sof15]. major [MP19a]. maler [Sta11c]. malerier [Sta11c]. Man [Cha18, Ekm16]. manned [Har11]. Manuel [CWR17]. map [Aal12, Pis14]. mapmaker [BN15, Nat16]. mapping [Aal15]. mascon [Ant22]. Maskelyne [Sma18]. Maurycy [JK16]. maximum [CW14]. Measurements [BSB^{+}20, AH13, Row23, San17, Sch15, WR18]. medieval [Pis14]. Mehmet [Mul24]. Melting [Car10, Riv11]. meridian [Mar16]. meteor [Spe21]. Meteorologica [Win23]. Meteorological [Nev14, Sta11c, Win23, dVKR^{+}23]. Meteorologiske [Sta11c]. meteorology [Mad23]. Milankovitch [Ate22, Gaj19]. millennia [ZC18]. mode [CI12]. model [She17]. modelling [Ant22]. models [Ate22, MKSG10]. modern [IR17, MKSG10, Ozc20]. Moltke [Sil12, Sta11c]. Moltkes [Sta11c]. Moos [GDN15]. mortality [HDPG^{+}13]. Mountains [Mei17]. Munich [Sof15]. my [Aka15, Lan24].

Nanabhoy [GDN15]. Natural [Ric07, Tay10]. naturalist [NH12]. navy [CH20]. necessary [FS12]. Netherlands [Car14, Ert10]. network [Win23]. Neumayer [SWS10]. newly [SN21]. Niemegk [Lin23a, Lin23b, Lin24]. noctilucent [DPR12]. nordlysets [Sta11c]. Northern [Bur16, Rai20]. Northwest [Med13]. northwestern [Aal15]. Norway [Hol12, Pet16]. Notes [DPR12, OO14, Mei17]. November [dVKR^{+}23].

Obituary [RS18, SG11]. observations [DPR12, Ege24, HR22, HS23, HR24, HVR23, Keu21, NT24, Pet16]. observatories [Apl20, ACFJ13, Lin23a, Lin23b, Lin24]. Observatory [Apl19a, BSB^{+}20, Bös21, TPCNA24, GDN15, PA12, Wei16, Bre21, Bur16, FN22, HR22, HR24, HVR23, Kot18, Mat13, NT24, Sof15, WFBM19]. observed [CAOV18]. occasion [IR17]. Ocean [Mil09, Sch13, SWGM^{+}19]. oldest [HGTJ21]. Oliva [DC18]. one [GDN15, HGTJ21]. Operation [Wan22]. organization [Win23]. Origin [She17]. origins [SL19]. other [GDN15]. Ottoman [OO14]. Our [Mil09, Dee11, MP19a, Sch13]. outstanding [MS12]. overview [Wit20]. Oxford [Apl19b]. Ozan [Mul24].

P [Jør11, Sta11b]. Painter [Sil12, Sta11c]. paintings [Sta11c]. Palatina [Win23]. Papers [Ric12]. Paradigm [Aka15]. Paris [ML16]. Part [Wan22, Hul11, Hae16, Hol12, Hul11, Lin23a, Lin23b, Lin24, MFN23, Oks11]. Partenkirchen [HS23]. participation [MFN23]. past [AFH^{+}19, Bös21, KSDJ19]. paved [BGK^{+}13]. peak [Ert11]. Peck [LGGAGL19]. pendulum [Agn20]. People [ZW21]. period [ZW21]. periploi [Shc18]. personal [Aka15, Bar15]. perspective [Bar15, Kot18]. Peter [MS12]. petroleum [Mul24]. photography [NT24]. Physical [Ozc20]. physicist [Kra13]. physics [Aka15, And21, FS17]. picture [Pis14]. Pioneer [Cha18, EB12, Aal15, CI12, Ekm16, FS17, GMdSM21, MS12, SL14, Spe21, Nag13]. pioneering [EB10, Mad23, NH12]. Pioneers [Ate22]. Pius [JK16]. Planet [Mil09, Sch13, MP19a]. Poland [JSBWL16, KSDJ19]. Pontus [Shc20]. possibilities [Puz19]. Potsdam [Lin23a, Lin23b, Lin24]. practices [CKC^{+}22]. Prague [HVR23]. pre [Ozc20]. pre-modern [Ozc20]. precision [Mar11]. prediction [KSDJ19, Woo20]. Preface [JIZ19b, Wil18]. present [AFH^{+}19, Bös21, Edw16, KSDJ19, MP19b]. prime [Mar16]. Processes [Fro10]. production [HS21]. professionals [Keu21]. project [Puz19]. promotion [SL19]. pseudoscientific [Edw16, Sud14]. Ptolemy [Mar11, Mar12, Mar16, San17, Shc18, Shc20]. pyramids [Sma18].

quality [Car21]. Quaternary [NH12]. quo [CKC^{+}22].

Radar [MP19b, Mat13]. Radars

[MH13, EZ13, PB11]. radiation [FS12]. Radio [WFBM19]. Radionuclides [BMvRS12, Fro10, Wil15]. rainfall [Ert11]. rainfall-runoff [Ert11]. Ralph [LGGAGL19]. Range [Thr18]. Rappaport [Ric12]. Rawer [RS18]. realization [Sma18]. reanalysis [dVKR^{+}23]. reception [Win23]. record [CWR17, DC18, Rai20]. recording [SBLW14]. Rectification [Mar12]. reference [Hul11]. Regener [CW14]. regional [Ant22, Wil18, Dam18, HWD^{+}18]. regularities [FGFL12]. Reinhold [HS23]. Reiter [HS23]. relations [Ert11]. Remains [Car14, Ert10]. Renaissance [Vaq17]. Reply [Sta11b]. Republic [ZW21, HS21]. research [Car21, CI12, Bur16, Har11, HVR23, JSBWL16, LASJ22, Med13, MS12, Nis10, SL14, SWS10, Spe21]. results [RSD^{+}13]. retrospective [FVZF14]. Review [BJB11, Car14, Nag13, Nat16, Ric12, Riv11, Sch13, Sil12, Tay10, Wil15, Ant22, Apl19a, Apl19b, Bre16, Cha18, Lan24, Liu24, Pet16]. revised [GT14]. revisions [Aal12]. Rhoda [Ric12]. Riccioli [San17]. ring [Bur12]. rivers [Ert11]. Roald [Dee11]. Robert [IR17, CI12]. Rock [Cat11]. Rocket [Thr18]. role [FS12, IZ16, LASJ22, MP19a, ZC18]. Rouelle [Ric12]. Royal [TPCNA24]. Rudzki [JK16]. Rügen [Wei16]. runoff [Ert11]. RWC [HWD^{+}18]. RWC-China [HWD^{+}18].

SABRE [NS14]. Saint [BGK^{+}13]. San [SBLW14]. sank [dVKR^{+}23]. Santin [BGK^{+}13]. satisfactory [Ert11]. scale [Mad23]. Scatter [MP19b, BGK^{+}13, WFBM19]. scholar [SS11]. School [JSBWL16]. Schröder [SG11]. Sci [Jør11, Sta11b]. Science [Apl19a, Mil09, BJB11, Bar15, Her10, Jør11, Kri12, Sta11a, Sta11b, Val17, Liu24, Sch13]. sciences [CKC^{+}22, MV19, Ozc20, AFH^{+}19, ZW21]. Scientific [HGTJ21, MFN23, Apl18, Car21, Edw16, IZ16, Sud14, Liu24]. Scientifique [Med13]. Scientist [Bre16, BJB11, Gla20]. scientists [Jac22, Keu21]. sea [Rai20, Row23]. Secchi [PA12]. Seddin [Lin23a, Lin23b, Lin24]. seismologist [Jac22]. seismology [Edw16, SL19]. Selection [Ric12]. Service [Len21, KSDJ19]. Shackleton [dVKR^{+}23]. Shadow [Car10, Riv11]. shaping [Bar15]. Sheeley [She17]. short [Bur12, Mat13]. Siegfried [Har11]. since [Ant22, Apl19b, HVR23, Med13, Sch15]. sites [HGTJ21]. Sixtieth [LGS^{+}20]. Skempton [LGGAGLdA19]. Skip [Apl18]. Slovakia [HGTJ21]. slow [Edw16]. small [Cas19]. Smolník [HGTJ21]. SMS [Gou22]. Societas [Win23]. Society [Car10, Riv11, Jac22, Gla20, Wit20]. Sodankylä [Bös21, NT24]. solar [Aka15, She17]. some [ML16, SN21]. Sondrestrom [MH13]. Sophus [MS12]. sources [Har20]. South [Aal15]. southern [Kot18]. Soviet [LGS^{+}20]. Space [BMvRS12, Jør11, Lan24, Liu24, Sta11b, Wil15, WKC18, Edw16, Har11, Nis10]. Spain [DC18, Vaq17, CAOV18]. spar [Kri12]. special [Apl20, Hul11, JIZ19b]. spectral [Sch15]. sphere [DÁ19]. spheres [JIZ19b]. Spiral [FGFL12]. St. [CWR17]. STARE [NS14]. STARE/SABRE [NS14]. States [Aal12]. statistical [San17]. Stauning [Jør11]. Stephen [Apl19b]. Steponas [Val17]. Stoffregen [SL14]. Størmer [Nag13, EB12]. story [NS14]. strain [Car21]. stratigraphic [CKC^{+}22]. structures [FGFL12]. Struve [Puz19]. Studies [RTR11, LASJ22, San17, Ric12]. Study [Mil09, Sch13]. Subdivision [Mei17]. subjected [DÁ19]. subsequent [BJB11]. sudden [SN21]. Sudiro [Edw16]. suggestion [Sma18]. Sun [Dee11]. Sungurlu [Mul24]. sunspots [HCB^{+}10]. supplementary [Mei17]. survey [EB19, RSD^{+}13]. Svante [Kra13]. Swedish [Hul11]. system [BGK^{+}13, Wan22]. Széchenyi [BSB^{+}20].

tales [SWGM^{+}19]. teacher [Gla20].

technical [Puz19]. technique [WFBM19]. Tegetmeyer [FS12]. tephra [LASJ22]. Tephrochronology [LASJ22]. Terrestrial [BMvRS12, Fro10, GT14, Wil15, Aka15]. text [GT14, Mei17]. their [CKC^{+}22, Gou22, HVR23]. Theodor [NH12]. theorist [Kra13]. Theory [BMvRS12, GT14, Wil15, Edw16, Sud14]. thinking [CKC^{+}22]. threads [And21]. thunder [Sch24]. Tidal [CH20, Woo20, WR18]. Tide [Woo20, Agn20, CWR17]. tides [Agn20]. Tim [Apl19b]. Time [Agn20, Ric07, Mad23, Tay10]. Timers [Fro10]. times [Ozc20]. Tohoku [SBLW14]. tourism [JSBWL16, JSBWL16]. Tracers [Fro10]. Tradition [HS21, Pis14]. traditions [And21]. trained [Keu21]. transition [Edw16, Sud14, ZC18]. transitions [Aka15]. translation [GT14, Mei17]. transtridecadal [HCB^{+}10]. Traute [HDPG^{+}13]. triangulation [Puz19]. Trieste [Rai20]. Tromholt [MS12]. Tromsø [Bur16, RS22]. tropical [Mad23]. Turkish [Mul24]. twentieth [GALGGL18, LGGAGL19]. two [ZC18].

Ukraine [EZ13]. ultraviolet [Sch15]. Understanding [dVKR^{+}23, MP19a]. unifying [ZC18]. Union [JIZ19b, LGS^{+}20]. unions [IZ16]. United [Aal12]. University [JSBWL16]. unmanned [Har11]. until [Dam18]. using [dVKR^{+}23].

vadis [CKC^{+}22]. variations [FGFL12]. Variscan [Mei17]. Vaupell [NH12]. vibrant [Cas19]. Victorian [Apl19a]. view [Aka15, Jac22]. virtual [Sma18]. voyage [WR18]. voyages [Gou22].

Wang [She17]. Warning [Dam18, HWD^{+}18, Wil18]. Warren [Liu24]. was [CH20]. way [BGK^{+}13]. Weather [Apl19b, HCB^{+}10, Nis10]. western [Mar16]. Westley [LGGAGLdA19]. whistler [CI12]. whistler-mode [CI12]. Wikipedia [Sch24]. Wilfried [SG11]. Willy [SL14]. wind [She17]. work [BN15]. works [San17]. world [Pis14, Ric12].

Year [LGS^{+}20]. years [Hae16, LASJ22, ML16].

Zafra [Vaq17]. Zealand [Row23]. Zeppelin [RSD^{+}13]. zero [ML16].

References

Aplin:2010:IE

[AAE^{+}10] K. Aplin, K. Arora, M. Ertsen, G. A. Good, G. P. Gregori, J. M. Herndon, T. Kikuchi, T. Kutsukake, T. V. Kuznetsova, S. Papamarinopoulos, R. J. Pellinen, P. Richet, K. Schlegel, W. Schröder, and V. Schwach. Introductory editorial. History of Geo- and Space Sciences, 1(1):1–2, ???? 2010. CODEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

//www.hist-geo-space-
sci.net/1/1/2010/.

Aalto:2012:EJJ

[Aal12] K. R. Aalto. Edwin James’ and John Hinton’s revisions of Maclure’s geologic map of the United States. History of Geoand Space Sciences, 3(1): 75–86, ???? 2012. CO-DEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

//www.hist-geo-space-
sci.net/3/75/2012/.

Aalto:2015:HKP

[Aal15]  K. R. Aalto.  Her-

mann Karsten, pioneer of geologic mapping in northwestern South America. History of Geo- and Space Sciences, 6(1):57– 63, ???? 2015. CO-DEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

//www.hist-geo-space-
sci.net/6/57/2015/.

Arora:2013:IHG

[ACFJ13] K. Arora, D. Cole, J. Urrutia Fucugauchi, and M. G. Johnson. Introduction: “History of geophysical institutes and observatories”. History of Geoand Space Sciences, 4(2): 71, ???? 2013. CO-DEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

//www.hist-geo-space-

sci.net/4/71/2013/hgss-
4-71-2013.pdf.

Allison:2019:IPP

[AFH^{+}19] Ian Allison, Charles Fierz, Regine Hock, Andrew Mackintosh, Georg Kaser, and Samuel U. Nussbaumer. IACS: past, present, and future of the International Association of Cryospheric Sciences. History of Geoand Space Sciences, 10(1): 97–107, ???? 2019. CO-DEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

//www.hist-geo-space-
sci.net/10/97/2019/.

Agnew:2020:TTP

[Agn20] Duncan C. Agnew. Time and tide: pendulum clocks and gravity tides. History of Geo- and Space Sciences, 11(2):215–224, ???? 2020. CODEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:/

/hgss.copernicus.org/
articles/11/215/2020/.

Aplin:2013:LKA

[AH13] K. L. Aplin and R. G. Harrison. Lord Kelvin’s atmospheric electricity measurements. History of Geoand Space Sciences, 4(2): 83–95, ???? 2013. CO-DEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

//www.hist-geo-space-
sci.net/4/83/2013/.

Akasofu:2015:PTS

[Aka15] S.-I. Akasofu. Paradigm transitions in solar–terrestrial physics from 1900: my personal view. History of Geoand Space Sciences, 6(1): 23–43, ???? 2015. CO-DEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

//www.hist-geo-space-
sci.net/6/23/2015/.

Anduaga:2021:FIP

[And21]  Aitor Anduaga.  The

formation of ionospheric physics — confluence of traditions and threads of continuity. History of Geoand Space Sciences, 12(1): 57–75, ???? 2021. CO-DEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

//hgss.copernicus.org/
articles/12/57/2021/.

Antoni:2022:RDM

[Ant22] Markus Antoni. A review of different mascon approaches for regional gravity field modelling since 1968. History of Geo- and Space Sciences, 13(2):205– 217, ???? 2022. CO-DEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

//hgss.copernicus.org/
articles/13/205/2022/.

Aplin:2018:AED

[Apl18]  Karen L. Aplin.  At-

mospheric electricity at Durham: the scientific contributions and legacy of J. A. (“Skip”) Chalmers (1904–1967). History of Geo- and Space Sciences, 9 (1):25–35, ???? 2018. CO-DEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

//www.hist-geo-space-
sci.net/9/25/2018/.

Aplin:2019:BRK

[Apl19a] Karen L. Aplin. Book review: Kew Observatory and the Evolution of Victorian Science 1840– 1910. History of Geoand Space Sciences, 10(1): 1–2, ???? 2019. CO-DEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

//www.hist-geo-space-
sci.net/10/1/2019/.

Aplin:2019:BRO

[Apl19b] Karen L. Aplin. Book review: Oxford Weather and Climate since 1767 by Stephen Burt and Tim Burt. History of Geo- and Space Sciences, 10(2):267– 268, ???? 2019. CO-DEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

//www.hist-geo-space-
sci.net/10/267/2019/.

Aplin:2020:ISI

[Apl20] Karen L. Aplin. Introduction to the special issue “Atmospheric electrical observatories”. History of Geo- and Space Sciences, 11(2):137–138, ???? 2020. CODEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

                                        //hgss.copernicus.org/
                                        articles/11/137/2020/.

Ates:2022:PIA

[Ate22] M. Efe Ates. Pioneers of the ice age models: a brief history from Agassiz to Milankovitch. History of Geoand Space Sciences, 13(1): 23–37, ???? 2022. CO-DEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

                                        //hgss.copernicus.org/
                                        articles/13/23/2022/.

Barry:2015:SCS

[Bar15] R. G. Barry. The shaping of climate science: half a century in personal perspective. History of Geoand Space Sciences, 6(2): 87–105, ???? 2015. CO-DEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

                                        //www.hist-geo-space-
                                        sci.net/6/87/2015/.

Bauer:2013:HSS

[BGK^{+}13] P. Bauer, A. Giraud, W. Kofman, M. Petit, and P. Waldteufel.

How the Saint Santin incoherent scatter system paved the way for a French involvement in EIS-CAT. History of Geoand Space Sciences, 4(2): 97–103, ???? 2013. CO-DEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

//www.hist-geo-space-
sci.net/4/97/2013/.

Barry:2011:RAB

[BJB11] R. G. Barry, J. Jania, and K. Birkenmajer. Review article: “A. B. Dobrowolski — the first cryospheric scientist — and the subsequent development of cryospheric science”. History of Geoand Space Sciences, 2(1): 75–79, ???? 2011. CO-DEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

//www.hist-geo-space-
sci.net/2/75/2011/.

Beer:2012:CRT

[BMvRS12] Jürg Beer, K. G. (Kenneth G.) McCracken, and R. von (Rudolf) Steiger. Cosmogenic Radionuclides: Theory and Applications in the Terrestrial and Space Environments. Physics of earth and space environments. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2012. ISBN 3-642-14650-3 (hardcover), 3-642-14651-1 (ebook). ISSN 1610-1677 (print), 1865-0678 (electronic). xvi + 426 pp. LCCN QC913 .B44 2012.

Brednich:2015:AKM

[BN15] Rolf Wilhelm Brednich and Sascha Nolden. Augustus Koch — mapmaker: the life and work of Augustus Koch (1834–1901): artist, designer, draughtsman and cartographer. Steele Roberts Aotearoa, Wellington, Aotearoa, New Zealand, 2015. ISBN 1-927242-87-8. 123 pp. LCCN GA1767.7.

Bosinger:2021:GOS

[Bös21] Tilmann Bösinger. The geophysical observatory in Sodankylä, Finland — past and present. History of Geo- and Space Sciences, 12(2):115–130, ???? 2021. CODEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

                                        //hgss.copernicus.org/
                                        articles/12/115/2021/.

Brednich:2016:BRJ

[Bre16] Rolf W. Brednich. Book review: James Hector. Explorer, Scientist, Leader. History of Geo- and Space Sciences, 7(1):65–66, ???? 2016. CODEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL

                                        https://www.hist-geo-

space-sci.net/7/65/2016/

.

Brekke:2021:LHO

[Bre21] Asgeir Brekke. The legacy of the Haldde Observatory. History of Geoand Space Sciences, 12(1): 1–9, ???? 2021. CO-DEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

//hgss.copernicus.org/
articles/12/1/2021/.

Bor:2020:MAE

[BSB^{+}20] József Bór, Gabriella Sátori, Veronika Barta, Karolina Szabóné-André, Judit Szendrői, Viktor Wesztergom, Tamás Bozóki, Attila Buzás, and Dávid Koronczay. Measurements of atmospheric electricity in the Széchenyi István Geophysical Observatory, Hungary. History of Geoand Space Sciences, 11(1): 53–70, ???? 2020. CO-DEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

//www.hist-geo-space-
sci.net/11/53/2020/.

Egeland:2012:RCS

[Bur12] A. Egeland W. J. Burke.

The ring current: a short biography. History of Geoand Space Sciences, 3(2): 131–142, ???? 2012. CO-DEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

                                        //www.hist-geo-space-
                                        sci.net/3/131/2012/.

Egeland:2016:ART

[Bur16] Alv Egeland William J. Burke. Auroral research at the Tromsø Northern Lights Observatory: the Harang directorship, 1928– 1946. History of Geoand Space Sciences, 7(1): 53–61, ???? 2016. CO-DEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

                                        //www.hist-geo-space-
                                        sci.net/7/53/2016/.

Carrasco:2018:GAJ

[CAOV18] Víctor M. S. Carrasco, Enric Aragonès, Jorge Ordaz, and José M. Vaquero. The Great Aurora of January 1770 observed in Spain. History of Geo- and Space Sciences, 9(2):133– 139, October 2018. ISSN 2190-5010 (print), 2190-5029 (electronic).

Carey:2010:SMG

[Car10] Mark (Mark P.) Carey. In the Shadow of Melting Glaciers: Climate Change and Andean Society. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2010. ISBN 0-19-539606-5 (hardcover), 0-19-539607-3 (paperback). vii + 273 pp. LCCN QC988.P4 C37 2010. URL http://

www.h-net.org/reviews/
showrev.php?id=32303.

Carey:2014:BRL

[Car14]  M. Carey.  Book re-

view: Locales of Happiness: Colonial Irrigation in the Netherlands East Indies and its Remains, 1830–1980. History of Geoand Space Sciences, 5(1): 73–74, ???? 2014. CO-DEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

//www.hist-geo-space-
sci.net/5/73/2014/hgss-
5-73-2014.pdf.

Carbonell:2021:HSQ

[Car21] Pablo J. Torres Carbonell.

On the high scientific quality of early research on strain and deformation fabrics (1835–1908). History of Geo- and Space Sciences, 12(2):197–216, ???? 2021. CODEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:/

/hgss.copernicus.org/
articles/12/197/2021/.

Cas:2019:ISB

[Cas19] Raymond A. F. Cas. IAV-CEI: from small beginnings to a vibrant international association. History of Geo- and Space Sciences, 10(1):181–191, ???? 2019. CODEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

//www.hist-geo-space-
sci.net/10/181/2019/.

Cathcart:2011:ARB

[Cat11]  R. B. Cathcart.  An-

thropic rock: a brief history. History of Geo- and Space Sciences, 2(1):57– 74, ???? 2011. CO-DEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

//www.hist-geo-space-
sci.net/2/57/2011/.

Carlsson-Hyslop:2020:HLT

[CH20]  Anna  Carlsson-Hyslop.

How the Liverpool Tidal Institute was established: industry, navy and academia. History of Geo- and Space Sciences, 11(2):139–156, ???? 2020. CODEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:/

/hgss.copernicus.org/
articles/11/139/2020/.

Chapman:2018:BRM

[Cha18] David S. Chapman. Book review: The Man behind Degree Celsius: A Pioneer in Investigating the Earth and its Changes. History of Geo- and Space Sciences, 9 (1):39–40, ???? 2018. CO-DEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

//www.hist-geo-space-
sci.net/9/39/2018/.

Carpenter:2012:RHP

[CI12] D. L. Carpenter and U. S.

Inan. Robert Helliwell, pioneer of whistler-mode research. History of Geoand Space Sciences, 3(1): 73–74, ???? 2012. CO-DEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

//www.hist-geo-space-
sci.net/3/73/2012/.

CarnierFragoso:2022:CES

[CKC^{+}22] Daniel Galvão Carnier Fragoso, Matheus Kuchenbecker, Antonio Jorge Campos Magalhães, Claiton Marlon Dos Santos Scherer, Guilherme Pederneiras Raja Gabaglia, and André Strasser. Cyclicity in Earth sciences, quo vadis? Essay on cycle concepts in geological thinking and their historical influence on stratigraphic practices. History of Geoand Space Sciences, 13(1): 39–69, ???? 2022. CO-DEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

//hgss.copernicus.org/
articles/13/39/2022/.

Carlson:2014:ERI

[CW14] P. Carlson and A. A. Watson. Erich Regener and the ionisation maximum of the atmosphere. History of Geo- and Space Sciences, 5(2):175–182, ???? 2014.

CODEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

                                        //www.hist-geo-space-
                                        sci.net/5/175/2014/.

Cartwright:2017:MJT

[CWR17] David E. Cartwright, Philip L. Woodworth, and Richard D. Ray. Manuel Johnson’s tide record at St. Helena. History of Geoand Space Sciences, 8(1): 9–19, ???? 2017. CO-DEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

                                        //www.hist-geo-space-
                                        sci.net/8/9/2017/.

Drewes:2019:IAG

[DÁ19] Hermann Drewes and József Ádám. The International Association of Geodesy: from an ideal sphere to an irregular body subjected to global change. History of Geo- and Space Sciences, 10(1):151–161, ???? 2019. CODEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

                                        //www.hist-geo-space-
                                        sci.net/10/151/2019/.

Damboldt:2018:RWC

[Dam18] Thomas Damboldt. The Regional Warning Center Darmstadt (from the 1960s until 1993). History of Geo- and Space Sciences, 9 (1):49–51, ???? 2018. CO-DEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

//www.hist-geo-space-
sci.net/9/49/2018/.

Dominguez-Castro:2018:ERB

[DC18] Fernando Dom’ınguez-Castro.

An early record of ball lightning: Oliva (Spain), 1619. History of Geoand Space Sciences, 9(1): 79–83, ???? 2018. CO-DEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

//www.hist-geo-space-
sci.net/9/79/2018/.

Egeland:2011:RAC

[Dee11] A. Egeland C. S. Deehr.

Roald Amundsen’s contributions to our knowledge of the magnetic fields of the Earth and the Sun. History of Geo- and Space Sciences, 2(2):99– 112, ???? 2011. CO-DEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

//www.hist-geo-space-
sci.net/2/99/2011/.

Dalin:2012:NHA

[DPR12] P. Dalin, N. Pertsev, and V. Romejko. Notes on historical aspects on the earliest known observations of noctilucent clouds. History of Geoand Space Sciences, 3(1): 87–97, ???? 2012. CO-DEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

                                        //www.hist-geo-space-
                                        sci.net/3/87/2012/.

deVos:2023:UDS

[dVKR^{+}23] Marc de Vos, Panagiotis Kountouris, Lasse Rabenstein, John Shears, Mira Suhrhoff, and Christian Katlein. Understanding the drift of Shackleton’s Endurance during its last days before it sank in November 1915, using meteorological reanalysis data. History of Geoand Space Sciences, 14(1): 1–13, ???? 2023. CO-DEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

                                        //hgss.copernicus.org/
                                        articles/14/1/2023/.

Egeland:2010:KBP

[EB10] A. Egeland and W. J. Burke. Kristian Birkeland’s pioneering investigations of geomagnetic disturbances. History of Geoand Space Sciences, 1(1): 13–24, ???? 2010. CO-DEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

                                        //www.hist-geo-space-
                                        sci.net/1/13/2010/.

Egeland:2012:CSA

[EB12] Alv Egeland and William J. Burke. Carl Størmer: Auroral Pioneer, volume 393 of Astrophysics and Space Science Library. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2012. ISBN 3-642-31456-2 (hardcover), 3-642-31457-0 (e-book). viii + 195 + 131 pp. LCCN QC971 .E34 2013. URL

http://link.springer.
com/book/10.1007/978-
3-642-31457-5;   http:
//public.eblib.com/choice/
publicfullrecord.aspx?
p=1030573;   http://site.
ebrary.com/id/10617872.

Egeland:2019:AHE

[EB19] Alv Egeland and William J.

Burke. Auroral hydrogen emissions: a historic survey. History of Geo- and Space Sciences, 10(1):201– 213, ???? 2019. CO-DEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

//www.hist-geo-space-
sci.net/10/201/2019/.

Edwards:2016:ISG

[Edw16] Matthew R. Edwards.

Indications from space geodesy, gravimetry and seismology for slow Earth expansion at present — comment on “The Earth expansion theory and its transition from scientific hypothesis to pseudoscientific belief” by Sudiro (2014). History of Geoand Space Sciences, 7(2):

125–133, ???? 2016. CO-DEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

                                        //www.hist-geo-space-
                                        sci.net/7/125/2016/. See

[Sud14].

Egeland:2024:CAO

[Ege24] Alv Egeland. Conjugate aurora observations by the Gjøa and Discovery expeditions. History of Geoand Space Sciences, 15(2): 27–39, ???? 2024. CO-DEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

                                        //hgss.copernicus.org/
                                        articles/15/27/2024/.

Ekman:2016:MBD

[Ekm16] Martin Ekman. The man behind “Degrees Celsius”: a pioneer in investigating the Earth and its changes. Summer Institute for Historical Geophysics, Godby, Åland Islands, Finland, 2016. ISBN 952-93-7732-0. 159 pp. LCCN ????

Ertsen:2010:LHC

[Ert10] Maurits W. Ertsen. Locales of Happiness Colonial Irrigation in the Netherlands East Indies and Its Remains, 1830–1980. VSSD, Delft, The Netherlands, 2010. ISBN 90-6562-241-1. ix + 238 pp. LCCN ????

Ertsen:2011:CSA

[Ert11]  M. W. Ertsen.  “A

not completely satisfactory attempt” — peak discharges and rainfallrunoff relations for Javanese rivers between 1880 and 1940. History of Geoand Space Sciences, 2(1): 39–55, ???? 2011. CO-DEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

//www.hist-geo-space-
sci.net/2/39/2011/.

Emelyanov:2013:HDR

[EZ13] L. Ya. Emelyanov and T. G. Zhivolup. History of the development of IS radars and founding of the Institute of Ionosphere in Ukraine. History of Geoand Space Sciences, 4(1): 7–17, ???? 2013. CO-DEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

//www.hist-geo-space-
sci.net/4/7/2013/.

Feldstein:2012:SSR

[FGFL12] Y. I. Feldstein, L. I. Gromova, M. Förster, and A. E. Levitin. Spiral structures and regularities in magnetic field variations and auroras. History of Geo- and Space Sciences, 3(1):1–31, ???? 2012. CO-DEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

                                        //www.hist-geo-space-
                                        sci.net/3/1/2012/.

Fujii:2022:HKM

[FN22] Ikuko Fujii and Shingo Nagamachi. History of Kakioka Magnetic Observatory. History of Geoand Space Sciences, 13(2): 147–170, ???? 2022. CO-DEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

                                        //hgss.copernicus.org/
                                        articles/13/147/2022/.

Froehlich:2010:ERH

[Fro10] Klaus F. O. Froehlich, editor. Environmental Radionuclides: Tracers and Timers of Terrestrial Processes, volume 16 of Radioactivity in the environment. Elsevier, Amsterdam, The Netherlands, 2010. ISBN 0-08-043873-3 (hardcover), 0-08-091329-6 (e-book). ISSN 1569-4860. xiii + 438 pp. LCCN QC795.8.R3 E58 2010; TK9400 .A63 2008. URL http://

                                        www.sciencedirect.com/
                                        science/bookseries/15694860/
                                        16.

Fricke:2012:ADC

[FS12] R. G. A. Fricke and K. Schlegel. 100th anniversary of the discovery of cosmic radiation: the role of Günther and Tegetmeyer in the development of the necessary instrumentation. History of Geoand Space Sciences, 3(2): 151–158, ???? 2012. CO-DEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

//www.hist-geo-space-
sci.net/3/151/2012/.

Fricke:2017:JEH

[FS17] Rudolf G. A. Fricke and Kristian Schlegel. Julius Elster and Hans Geitel — dioscuri of physics and pioneer investigators in atmospheric electricity. History of Geo- and Space Sciences, 8(1):1–7, ???? 2017. CODEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

//www.hist-geo-space-
sci.net/8/1/2017/.

Feldstein:2014:IAL

[FVZF14] Y. I. Feldstein, V. G.

Vorobjev, V. L. Zverev, and M. Förster. Investigations of the auroral luminosity distribution and the dynamics of discrete auroral forms in a historical retrospective. History of Geoand Space Sciences, 5(1): 81–134, ???? 2014. CO-DEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

//www.hist-geo-space-
sci.net/5/81/2014/.

Gajic:2019:CCM

[Gaj19] Nenad Gajic. The curious case of the Milankovitch calendar. History of Geoand Space Sciences, 10(2): 235–243, ???? 2019. CO-DEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

//www.hist-geo-space-
sci.net/10/235/2019/.

Galindo-Aires:2018:CKE

[GALGGL18] Rubén Galindo-Aires, Antonio Lara-Galera, and Gonzalo Guillán-Llorente. Contribution to the knowledge of early geotechnics during the twentieth century: Arthur Casagrande. History of Geo- and Space Sciences, 9(2):107–123, ???? 2018. CODEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

//www.hist-geo-space-
sci.net/9/107/2018/.

Gawali:2015:CAM

[GDN15] P. B. Gawali, M. G. Doiphode, and R. N. Nimje. Colaba–Alibag magnetic observatory and Nanabhoy Moos: the influence of one over the other. History of Geoand Space Sciences, 6(2): 107–131, ???? 2015. CO-DEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

//www.hist-geo-space-
sci.net/6/107/2015/.

Glassmeier:2020:KFA

[Gla20] Karl-Heinz Glassmeier.

Karl Friedrich Almstedt — scientist, teacher, and co-founder of the German Geophysical Society. History of Geo- and Space Sciences, 11(1):71– 80, ???? 2020. CO-DEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

//www.hist-geo-space-
sci.net/11/71/2020/.

Guillan-Llorente:2025:CKE

[GLMMLGGA25] Gonzalo Guillán-Llorente, Belén Muñoz-Medina, Antonio Lara-Galera, and Rubén Galindo-Aires. Contribution to the knowledge of early geotechnics during the 20th century: Laurits Bjerrum. History of Geoand Space Sciences, 16(1): 1–12, ???? 2025. CO-DEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

//hgss.copernicus.org/
articles/16/1/2025/.

Gois-Marques:2021:BGH

[GMdSM21] Carlos A. Góis-Marques, Miguel Menezes de Sequeira, and José Madeira. The bicentenary of Georg Hartung, a German pioneer geologist, explorer, and illustrator. History of Geo- and Space Sciences, 12(2):217–223, ???? 2021. CODEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

                                        //hgss.copernicus.org/
                                        articles/12/217/2021/.

Gould:2022:HCS

[Gou22] W. John Gould. HMS Challenger and SMS Gazelle — their 19th century voyages compared. History of Geo- and Space Sciences, 13(2):171–204, ???? 2022. CODEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

                                        //hgss.copernicus.org/
                                        articles/13/171/2022/.

Glassmeier:2014:CFG

[GT14] K.-H. Glassmeier and B. T. Tsurutani. Carl Friedrich Gauss — general theory of terrestrial magnetism — a revised translation of the German text. History of Geo- and Space Sciences, 5 (1):11–62, ???? 2014. CO-DEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

                                        //www.hist-geo-space-
                                        sci.net/5/11/2014/.

Haerendel:2016:HEP

[Hae16] Gerhard Haerendel. History of EISCAT — Part 4: On the German contribution to the early years of EISCAT. History of Geoand Space Sciences, 7(2): 67–72, ???? 2016. CO-DEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

//www.hist-geo-space-
sci.net/7/67/2016/.

Haig:1983:JHE

[Hai83] Bruce Haig. James Hector explorer. Following historic trails. Alberta Historical Resources Foundation, Calgary, Alberta, Canada, 1983. ISBN 0-920490-35-2 (paperback). 51 pp. LCCN F1060.8.

Hartmann:2011:BUM

[Har11] G. K. Hartmann. “Between unmanned and manned space research”: Dr. Siegfried J. Bauer for his 80th birthday. History of Geo- and Space Sciences, 2 (1):81–82, ???? 2011. CO-DEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

//www.hist-geo-space-
sci.net/2/81/2011/.

Harrison:2020:BCC

[Har20] R. Giles Harrison. Behind the curve: a comparison of historical sources for the Carnegie curve of the global atmospheric electric circuit. History of Geoand Space Sciences, 11(2): 207–213, ???? 2020. CO-DEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

//hgss.copernicus.org/
articles/11/207/2020/.

Halberg:2010:EGT

[HCB^{+}10] F. Halberg, G. Cornélissen, K.-H. Bernhardt, M. Sampson, O. Schwartzkopff, and D. Sonntag. Egeson’s (George’s) transtridecadal weather cycling and sunspots. History of Geo- and Space Sciences, 1(2):49–61, ???? 2010. CODEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL

https://www.hist-geo-
space-sci.net/1/49/2010/

.

Halberg:2013:DCH

[HDPG^{+}13] F. Halberg, N. Düll-Pfaff, L. Gumarova, T. A. Zenchenko, O. Schwartzkopff, E. M. Freytag, J. Freytag, and G. Cornelissen. 27day cycles in human mortality: Traute and Bernhard Düll. History of Geoand Space Sciences, 4(1): 47–59, ???? 2013. CO-DEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

//www.hist-geo-space-
sci.net/4/47/2013/.

Herndon:2010:ISH

[Her10] J. M. Herndon. Inseparability of science history and discovery. History of Geoand Space Sciences, 1(1): 25–41, ???? 2010. CO-DEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

//www.hist-geo-space-
sci.net/1/25/2010/.

Hroncek:2021:SJO

[HGTJ21] Pavel Hroncek, Bohuslava Gregorová, Dana Tometzová, and Milos Jesenský. Scientific journeys to one of the oldest copper cementation sites in Central Europe (Smolník, Slovakia). History of Geoand Space Sciences, 12(2): 179–196, ???? 2021. CO-DEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

//hgss.copernicus.org/
articles/12/179/2021/.

Holt:2012:HEP

[Hol12] O. Holt. History of EIS-CAT — Part 3: The early history of EISCAT in Norway. History of Geo- and Space Sciences, 3(1):47– 52, ???? 2012. CO-DEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

//www.hist-geo-space-
sci.net/3/47/2012/.

Harrison:2022:AEO

[HR22] R. Giles Harrison and John C. Riddick. Atmospheric electricity observations at Lerwick Geophysical Observatory. History of Geo- and Space Sciences, 13(2):133–146, ???? 2022. CODEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

//hgss.copernicus.org/
articles/13/133/2022/.

Harrison:2024:AEO

[HR24] R. Giles Harrison and John C. Riddick. Atmospheric electricity observations at Eskdalemuir Geophysical Observatory. History of Geoand Space Sciences, 15(1): 5–16, ???? 2024. CO-DEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

//hgss.copernicus.org/
articles/15/5/2024/.

Hanek:2021:TGI

[HS21] Pavel Hánek and Pavel Hánek Sr. Tradition of geodetic instruments production in the Czech Republic. History of Geo- and Space Sciences, 12(2):171– 178, ???? 2021. CO-DEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

//hgss.copernicus.org/
articles/12/171/2021/.

Harrison:2023:AEO

[HS23] R. Giles Harrison and Kristian Schlegel. Atmospheric electricity observations by Reinhold Reiter around Garmisch-Partenkirchen. History of Geo- and Space Sciences, 14(1):71–75, ???? 2023. CODEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

//hgss.copernicus.org/
articles/14/71/2023/.

Hultqvist:2011:HEP

[Hul11] B. Hultqvist. History of EISCAT — Part 1: On the early history of EIS-CAT with special reference to the Swedish part of it. History of Geo- and Space Sciences, 2(2):115– 121, ???? 2011. CO-DEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

//www.hist-geo-space-
sci.net/2/115/2011/.

Hejda:2023:HGO

[HVR23] Pavel Hejda, Fridrich Valach, , and Milos Revallo. Historical geomagnetic observations from Prague Observatory (since 1839) and their contribution to geomagnetic research. History of Geoand Space Sciences, 14(1): 51–60, ???? 2023. CO-DEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

//hgss.copernicus.org/
articles/14/51/2023/.

He:2018:BHR

[HWD^{+}18] Han He, Huaning Wang, Zhanle Du, Xin Huang, Yan Yan, Xinghua Dai, Juan Guo, and Jialong Wang. A brief history of Regional Warning Center China (RWC-China). History of Geoand Space Sciences, 9(1): 41–47, ???? 2018. CO-DEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

                                        //www.hist-geo-space-
                                        sci.net/9/41/2018/.

Ihde:2017:FRH

[IR17] Johannes Ihde and Andreas Reinhold. Friedrich Robert Helmert, founder of modern geodesy, on the occasion of the centenary of his death. History of Geoand Space Sciences, 8(2): 79–95, ???? 2017. CO-DEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

                                        //www.hist-geo-space-
                                        sci.net/8/79/2017/.

Ismail-Zadeh:2016:GIR

[IZ16] Alik Ismail-Zadeh. Geoscience international: the role of scientific unions. History of Geo- and Space Sciences, 7(2):103–123, ???? 2016. CODEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

                                        //www.hist-geo-space-
                                        sci.net/7/103/2016/.

Ismail-Zadeh:2019:IBE

[IZJ19] Alik Ismail-Zadeh and Jo Ann Joselyn. IUGG: beginning, establishment, and early development (1919–1939). History of Geo- and Space Sciences, 10(1):25–44, ???? 2019. CODEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

//www.hist-geo-space-
sci.net/10/25/2019/.

Jacobsen:2022:IGI

[Jac22] Lif Lund Jacobsen. Intellectually gifted but inherently fragile — society’s view of female scientists as experienced by seismologist Inge Lehmann up to 1930. History of Geoand Space Sciences, 13(1): 83–92, ???? 2022. CO-DEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

//hgss.copernicus.org/
articles/13/83/2022/.

Joselyn:2019:IE

[JIZ19a] Jo Ann Joselyn and Alik Ismail-Zadeh. IUGG evolves (1940–2000). History of Geo- and Space Sciences, 10(1):45–72, ???? 2019. CODEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL

https://www.hist-geo-
space-sci.net/10/45/2019/

.

Joselyn:2019:PSI

[JIZ19b] Jo Ann Joselyn and Alik Ismail-Zadeh. Preface to the special issue “The International Union of Geodesy and Geophysics: from different spheres to a common globe”. History of Geo- and Space Sciences, 10(1):17–24, ???? 2019. CODEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

                                        //www.hist-geo-space-
                                        sci.net/10/17/2019/.

Joselyn:2019:IC

[JIZB^{+}19] Jo Ann Joselyn, Alik Ismail-Zadeh, Tom Beer, Harsh Gupta, Masaru Kono, Uri Shamir, Michael Sideris, and Kathryn Whaler. IUGG in the 21st century. History of Geoand Space Sciences, 10(1): 73–95, ???? 2019. CO-DEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

                                        //www.hist-geo-space-
                                        sci.net/10/73/2019/.

Jackowski:2016:MPR

[JK16] Antoni Jackowski and Kazimierz Krzemień. Maurycy Pius Rudzki and the birth of geophysics. History of Geo- and Space Sciences, 7 (1):23–25, ???? 2016. CO-DEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

                                        //www.hist-geo-space-
                                        sci.net/7/23/2016/.

Jorgensen:2011:CDA

[Jør11] T. S. Jørgensen. Comment on “Danish auroral science history” by P. Stauning in Hist. Geo Space Sci., 2, 1– 28, 2011. History of Geoand Space Sciences, 2(2):

85–86, ???? 2011. CO-DEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

//www.hist-geo-space-
sci.net/2/85/2011/. See

[Sta11a, Sta11b].

Jackowski:2016:GTR

[JSBWL16] Antoni Jackowski, Izabela Soljan, Elzbieta Bilska-Wodecka, and Justyna Liro. Geographical tourism research and education at the Jagiellonian University School of Tourism in Poland (1936–1939). History of Geo- and Space Sciences, 7(2):91–101, ???? 2016. CODEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL

https://www.hist-geo-
space-sci.net/7/91/2016/

.

Keul:2021:BHB

[Keu21] Alexander G. Keul. A brief history of ball lightning observations by scientists and trained professionals. History of Geoand Space Sciences, 12(1): 43–56, ???? 2021. CO-DEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

//hgss.copernicus.org/
articles/12/43/2021/.

Kotze:2018:HMO

[Kot18] Pieter B. Kotzé. Hermanus Magnetic Observatory: a historical perspective of geomagnetism in southern Africa. History of Geoand Space Sciences, 9(2): 125–131, ???? 2018. CO-DEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

//www.hist-geo-space-
sci.net/9/125/2018/.

Kragh:2013:SAC

[Kra13] Helge Kragh. Svante Arrhenius, cosmical physicist and auroral theorist. History of Geo- and Space Sciences, 4(2):61– 69, ???? 2013. CO-DEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

//www.hist-geo-space-
sci.net/4/61/2013/.

Kragh:2015:EED

[Kra15] Helge Kragh. Expanding Earth and declining gravity: a chapter in the recent history of geophysics. History of Geoand Space Sciences, 6(1): 45–55, ???? 2015. CO-DEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

//www.hist-geo-space-
sci.net/6/45/2015/.

Kristjansson:2012:ISL

[Kri12] L. Kristjánsson. Iceland spar and its legacy in science. History of Geo- and Space Sciences, 3(1):117– 126, ???? 2012. CO-DEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

//www.hist-geo-space-
sci.net/3/117/2012/.

Klos:2019:HPS

[KSDJ19] Zbigniew Klos, Iwona Stanislawska, and Beata Dziak-Jankowska. Heliogeophysical prediction service in Poland: past, present and future. History of Geo- and Space Sciences, 10(1):193–199, ???? 2019. CODEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL

https://www.hist-geo-
space-sci.net/10/193/
2019/.

Lanzerotti:2024:BRM

[Lan24] Louis J. Lanzerotti. Book review: My Life in Space Exploration, by G. Haerendel. History of Geo- and Space Sciences, 15(2):43– 44, ???? 2024. CO-DEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

//hgss.copernicus.org/
articles/15/43/2024/.

Lowe:2022:GTS

[LASJ22] David J. Lowe, Peter M.

Abbott, Takehiko Suzuki, and Britta J. L. Jensen. Global tephra studies: role and importance of the international tephra research group “Commission on tephrochronology” in its first 60 years. History of Geo- and Space Sciences, 13(2):93–132, ???? 2022. CODEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

                                        //hgss.copernicus.org/
                                        articles/13/93/2022/.

Lenhardt:2021:HGS

[Len21] Wolfgang A. Lenhardt. The history of the Geophysical Service of Austria. History of Geo- and Space Sciences, 12(1):11– 19, ???? 2021. CO-DEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

                                        //hgss.copernicus.org/
                                        articles/12/11/2021/.

Lara-Galera:2019:CKEa

[LGGAGL19] Antonio Lara-Galera, Rubén Galindo-Aires, and Gonzalo Guillán-Llorente. Contribution to the knowledge of early geotechnics during the twentieth century: Ralph Peck. History of Geo- and Space Sciences, 10(1):3–15, ???? 2019. CO-DEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

                                        //www.hist-geo-space-
                                        sci.net/10/3/2019/.

Lara-Galera:2019:CKEb

[LGGAGLdA19] Antonio Lara-Galera, Rubén Galindo-Aires, Gonzalo Guillán-Llorente, and Vicente Alcaraz Carrillo de Albornoz. Contribution to the knowledge of early geotechnics during the 20th century: Alec Westley Skempton. History of Geo- and Space Sciences, 10(2):225–234, ???? 2019. CODEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL

https://www.hist-geo-
space-sci.net/10/225/
2019/.

Lyubovtseva:2020:SAI

[LGS^{+}20] Yulia S. Lyubovtseva, Alexei D. Gvishiani, Anatoly A. Soloviev, Olga O. Samokhina, and Roman I. Krasnoperov. Sixtieth anniversary of the International Geophysical Year (1957–2017) — contribution of the Soviet Union. History of Geo- and Space Sciences, 11(2):157–171, ???? 2020. CODEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:/

/hgss.copernicus.org/
articles/11/157/2020/.

Linthe:2023:HPS

[Lin23a] Hans-Joachim Linthe. History of the Potsdam, Seddin and Niemegk geomagnetic observatories — Part 1: Potsdam. History of Geo- and Space Sciences, 14(1):23–31, ???? 2023. CODEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

                                        //hgss.copernicus.org/
                                        articles/14/23/2023/.

Linthe:2023:HPSb

[Lin23b] Hans-Joachim Linthe. History of the Potsdam, Seddin and Niemegk geomagnetic observatories — Part 2: Seddin. History of Geoand Space Sciences, 14(1): 43–50, ???? 2023. CO-DEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

                                        //hgss.copernicus.org/
                                        articles/14/43/2023/.

Linthe:2024:HPS

[Lin24] Hans-Joachim Linthe. History of the Potsdam, Seddin and Niemegk geomagnetic observatories — Part 3: Niemegk. History of Geo- and Space Sciences, 15(2):45–69, ???? 2024. CODEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

                                        //hgss.copernicus.org/
                                        articles/15/45/2024/.

Liu:2024:BRS

[Liu24] Huixin Liu. Book review: Scientific Debates in Space Science, by Warren David Cummings and Louis J. Lanzerotti. History of Geoand Space Sciences, 15(2): 41–42, ???? 2024. CO-DEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

                                        //hgss.copernicus.org/
                                        articles/15/41/2024/.

Madden:2023:PTD

[Mad23] Roland A. Madden. A pioneering time of discoveries in large-scale tropical meteorology: 1960 through 1972. History of Geoand Space Sciences, 14(1): 33–41, ???? 2023. CO-DEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

//hgss.copernicus.org/
articles/14/33/2023/.

Marx:2011:PPG

[Mar11] C. Marx. On the precision of Ptolemy’s geographic coordinates in his Geographike Hyphegesis. History of Geo- and Space Sciences, 2(1):29– 37, ???? 2011. CO-DEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

//www.hist-geo-space-
sci.net/2/29/2011/.

Marx:2012:RAG

[Mar12]  C. Marx.  Rectifica-

tion of the ancient geographic coordinates in Ptolemy’s Geographike Hyphegesis. History of Geoand Space Sciences, 3(1): 99–112, ???? 2012. CO-DEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

//www.hist-geo-space-
sci.net/3/99/2012/.

Marx:2016:WCA

[Mar16] Christian Marx. The western coast of Africa in Ptolemy’s Geography and the location of his prime meridian. History of Geoand Space Sciences, 7(1): 27–52, ???? 2016. CO-DEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

//www.hist-geo-space-
sci.net/7/27/2016/.

Mathews:2013:SHG

[Mat13]  J. D. Mathews.  A

short history of geophysical radar at Arecibo Observatory. History of Geoand Space Sciences, 4(1): 19–33, ???? 2013. CO-DEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

//www.hist-geo-space-
sci.net/4/19/2013/.

Medina:2013:CIS

[Med13]  F. Medina.  Contribu-

tion of the “Institut Scientifique Chérifien” to the development of geoscientific research in Northwest Africa since its creation in 1914. History of Geoand Space Sciences, 4(2): 73–82, ???? 2013. CO-DEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

//www.hist-geo-space-
sci.net/4/73/2013/.

Meinhold:2017:FKS

[Mei17] Guido Meinhold. Franz Kossmat — subdivision of the Variscan Mountains — a translation of the German text with supplementary notes. History of Geoand Space Sciences, 8(1): 29–51, ???? 2017. CO-DEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

//www.hist-geo-space-
sci.net/8/29/2017/.

Matuura:2023:HEP

[MFN23] Nobuo Matuura, Ryoichi Fujii, , and Satonori Nozawa. History of EIS-CAT — Part 6: the participation of Japan in the EISCAT Scientific Association. History of Geoand Space Sciences, 14(1): 61–69, ???? 2023. CO-DEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

//hgss.copernicus.org/
articles/14/61/2023/.

McCready:2013:CSR

[MH13] M. A. McCready and C. J. Heinselman. The Chatanika and Sondrestrom radars — a brief history. History of Geoand Space Sciences, 4(1): 1–6, ???? 2013. CO-DEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

//www.hist-geo-space-
sci.net/4/1/2013/.

Mills:2009:FEO

[Mil09]  Eric L. Mills.  The

Fluid Envelope of Our Planet: How the Study of Ocean Currents Became a Science. University of Toronto Press, Toronto, ON, Canada, 2009. ISBN 0-8020-9697-2. xii + 434 pp. LCCN GC29 .M55 2009.

Mandea:2010:AHC

[MKSG10] M. Mandea, M. Korte, A. Soloviev, and A. Gvishiani. Alexander von Humboldt’s charts of the Earth’s magnetic field: an assessment based on modern models. History of Geo- and Space Sciences, 1 (2):63–76, ???? 2010. CO-DEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

//www.hist-geo-space-
sci.net/1/63/2010/.

Mandea:2016:ASY

[ML16] Mioara Mandea and Jean-Louis Le Mouël. After some 350 years — zero declination again in Paris. History of Geoand Space Sciences, 7(2): 73–77, ???? 2016. CO-DEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

//www.hist-geo-space-
sci.net/7/73/2016/.

Mandea:2019:IMR

[MP19a] Mioara Mandea and Eduard Petrovský. IAGA: a major role in understanding our magnetic planet. History of Geo- and Space Sciences, 10(1):163–172, ???? 2019. CODEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

//www.hist-geo-space-
sci.net/10/163/2019/.

Medvedev:2019:IIS

[MP19b] Andrey V. Medvedev and Alexander P. Potekhin. Irkutsk incoherent scatter radar: history, present and future. History of Geoand Space Sciences, 10(2): 215–224, ???? 2019. CO-DEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

//www.hist-geo-space-
sci.net/10/215/2019/.

Moss:2012:SPT

[MS12] K. Moss and P. Stauning.

Sophus Peter Tromholt: an outstanding pioneer in auroral research. History of Geo- and Space Sciences, 3 (1):53–72, ???? 2012. CO-DEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

//www.hist-geo-space-
sci.net/3/53/2012/.

Mulayim:2024:MOS

[Mul24] Oguz Mulayim. Mehmet Ozan Sungurlu, the legendary Turkish petroleum geologist. History of Geoand Space Sciences, 15(1): 1–4, ???? 2024. CO-DEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

//hgss.copernicus.org/
articles/15/1/2024/.

MacCracken:2019:ICI

[MV19] Michael C. MacCracken and Hans Volkert. IAMAS: a century of international cooperation in atmospheric sciences. History of Geoand Space Sciences, 10(1): 119–136, ???? 2019. CO-DEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

//www.hist-geo-space-
sci.net/10/119/2019/.

Nagarajan:2013:BRC

[Nag13] N. Nagarajan. Book review: Carl Størmer, Auroral Pioneer. History of Geo- and Space Sciences, 4(2):107–108, ???? 2013. CODEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

//www.hist-geo-space-
sci.net/4/107/2013/hgss-
4-107-2013.pdf.

Nathan:2016:BRA

[Nat16] S. Nathan. Book review: Augustus Koch — Mapmaker. History of Geoand Space Sciences, 7(1):

63–64, ???? 2016. CO-DEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

//www.hist-geo-space-
sci.net/7/63/2016/.

Nevanlinna:2014:EHF

[Nev14]  H. Nevanlinna.  On

the early history of the Finnish Meteorological Institute. History of Geoand Space Sciences, 5(1): 75–80, ???? 2014. CO-DEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

//www.hist-geo-space-
sci.net/5/75/2014/.

Nielsen:2012:CTV

[NH12] J. K. Nielsen and S. Helama.

Christian Theodor Vaupell, a Danish 19th century naturalist and a pioneering developer of the Quaternary geoscience. History of Geo- and Space Sciences, 3(2):143–150, ???? 2012. CODEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

//www.hist-geo-space-
sci.net/3/143/2012/.

Nishida:2010:EJC

[Nis10] A. Nishida. Early Japanese contributions to space weather research (1945– 1960). History of Geoand Space Sciences, 1(1): 1–12, ???? 2010. CO-DEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

                                        //www.hist-geo-space-
                                        sci.net/1/1/2010/.

Nielsen:2014:SSS

[NS14] E. Nielsen and W. Schmidt. The STARE/SABRE story. History of Geo- and Space Sciences, 5(1):63–72, ???? 2014. CODEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL

                                        https://www.hist-geo-
                                        space-sci.net/5/63/2014/

.

Nevanlinna:2024:EAP

[NT24] Heikki Nevanlinna and Eija I. Tanskanen. Early auroral photography and observations at the Sodankylä Geophysical Observatory in Finland, 1927– 1929. History of Geoand Space Sciences, 15(1): 17–25, ???? 2024. CO-DEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

                                        //hgss.copernicus.org/
                                        articles/15/17/2024/.

Oksman:2011:HEP

[Oks11] J. Oksman. History of EISCAT — Part 2: The early history of EISCAT in Finland. History of Geoand Space Sciences, 2(2): 123–128, ???? 2011. CO-DEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

                                        //www.hist-geo-space-
                                        sci.net/2/123/2011/.

Ozcep:2014:NHG

[OO14] F. Ozcep and T. Ozcep.

Notes on the history of geophysics in the Ottoman Empire. History of Geoand Space Sciences, 5(2): 163–174, ???? 2014. CO-DEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

//www.hist-geo-space-
sci.net/5/163/2014/.

Ozcep:2020:PES

[Ozc20] Ferhat Ozcep. Physical earth and its sciences in Istanbul: a journey from premodern (Islamic) to modern times. History of Geoand Space Sciences, 11(2): 173–198, ???? 2020. CO-DEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

//hgss.copernicus.org/
articles/11/173/2020/.

Ptitsyna:2012:FSF

[PA12] N. Ptitsyna and A. Altamore. Father Secchi and the first Italian magnetic observatory. History of Geo- and Space Sciences, 3 (1):33–45, ???? 2012. CO-DEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

//www.hist-geo-space-
sci.net/3/33/2012/.

Pellinen:2011:IHI

[PB11] R. Pellinen and A. Brekke.

Introduction: “The history of ionospheric radars”. History of Geo- and Space Sciences, 2(2):113–114, ???? 2011. CODEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

//www.hist-geo-space-
sci.net/2/113/2011/hgss-
2-113-2011.pdf.

Pettersen:2016:HRG

[Pet16] Bjørn Ragnvald Pettersen. A historical review of gravimetric observations in Norway. History of Geoand Space Sciences, 7(2): 79–89, ???? 2016. CO-DEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

//www.hist-geo-space-
sci.net/7/79/2016/.

Pischke:2014:EMT

[Pis14] G. Pischke. The Ebstorf Map: tradition and contents of a medieval picture of the world. History of Geo- and Space Sciences, 5(2):155–161, ???? 2014. CODEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

//www.hist-geo-space-
sci.net/5/155/2014/.

Puziene:2019:SGA

[Puz19] Ruta Puziene. The Struve Geodetic Arc: the development of the triangulation, technical possibilities, and the initiation of the project. History of Geoand Space Sciences, 10(2): 269–277, ???? 2019. CO-DEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

//www.hist-geo-space-
sci.net/10/269/2019/.

Raicich:2020:SLR

[Rai20] Fabio Raicich. A 1782– 1794 sea level record at Trieste (northern Adriatic). History of Geoand Space Sciences, 11(1): 1–14, ???? 2020. CO-DEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

//www.hist-geo-space-
sci.net/11/1/2020/.

Richet:2007:NHT

[Ric07] Pascal Richet. A Natural History of Time. University of Chicago Press, Chicago, IL, USA, 2007. ISBN 0-226-71287-7 (hardcover). xiv + 471 pp. LCCN QE508 .R5413 2007. URL http:

//www.loc.gov/catdir/
enhancements/fy0707/2006033992-
b.html;   http://www.
loc.gov/catdir/enhancements/
fy0707/2006033992-d.html;
http://www.loc.gov/catdir/
toc/ecip073/2006033992.
html.

Richet:2012:BRH

[Ric12] P. Richet. Book review:

From Hooke and Leibniz to Rouelle and Lavoisier, the eventful world of early geologists. A review of “Studies on Eighteenth-Century Geology, a Selection of Papers by Rhoda Rappaport”. History of Geoand Space Sciences, 3(1): 113–115, ???? 2012. CO-DEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

//www.hist-geo-space-
sci.net/3/113/2012/hgss-
3-113-2012.pdf.

Rivera:2011:BRS

[Riv11]  A. Rivera.  Book re-

view: “In the Shadow of Melting Glaciers. Climate change and Andean Society”. History of Geoand Space Sciences, 2(2): 97–98, ???? 2011. CO-DEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

//www.hist-geo-space-
sci.net/2/97/2011/hgss-
2-97-2011.pdf.

Rowe:2023:NZF

[Row23]  Glen H. Rowe.  New

Zealand’s first gaugebased sea level measurements. History of Geoand Space Sciences, 14(1): 77–92, ???? 2023. CO-DEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

//hgss.copernicus.org/
articles/14/77/2023/.

Rosbjerg:2019:IBH

[RR19] Dan Rosbjerg and John Rodda. IAHS: a brief history of hydrology. History of Geo- and Space Sciences, 10(1):109–118, ???? 2019. CODEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL

https://www.hist-geo-
space-sci.net/10/109/
2019/.

Reinisch:2018:OKR

[RS18] Bodo W. Reinisch and Kristian Schlegel. Obituary: Karl Rawer (1913– 2018). History of Geoand Space Sciences, 9(1): 105–106, ???? 2018. CO-DEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

//www.hist-geo-space-
sci.net/9/105/2018/.

Rietveld:2022:HTI

[RS22] Michael T. Rietveld and Peter Stubbe. History of the Tromsø ionosphere heating facility. History of Geo- and Space Sciences, 13(1):71–82, ???? 2022. CODEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

//hgss.copernicus.org/
articles/13/71/2022/.

Raspopov:2013:FAS

[RSD^{+}13] O. M. Raspopov, S. N.

Sokolov, I. M. Demina, R. Pellinen, and A. A.

Petrova. The first aeromagnetic survey in the Arctic: results of the Graf Zeppelin airship flight of 1931. History of Geoand Space Sciences, 4(1): 35–46, ???? 2013. CO-DEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

//www.hist-geo-space-
sci.net/4/35/2013/.

Rappaport:2011:SEC

[RTR11] Rhoda Rappaport, Kenneth L. Taylor, and M. J. S. Rudwick, editors. Studies on Eighteenth-century Geology. Variorum collected studies series. Ashgate Variorum, Farnham, UK, 2011. ISBN 1-4094-2959-8 (hardcover). 350 (est.) pp. LCCN QE13.E85 R358 2011.

Santoro:2017:SAL

[San17] Luca Santoro. A statistical approach to latitude measurements: Ptolemy’s and Riccioli’s geographical works as case studies. History of Geo- and Space Sciences, 8(2):69– 77, ???? 2017. CO-DEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

//www.hist-geo-space-
sci.net/8/69/2017/.

Steffen:2014:SFT

[SBLW14] H. Steffen, W. Brunk, M. Leven, and U. Wedeken.

From San Francisco to Tohoku — 111 yr of continuous earthquake recording in Göttingen. History of Geo- and Space Sciences, 5 (1):1–10, ???? 2014. CO-DEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

//www.hist-geo-space-
sci.net/5/1/2014/.

Schroder:2010:DAJ

[Sch10] W. Schröder. The development of the aurora of 18 January 1770. History of Geo- and Space Sciences, 1 (1):45–48, ???? 2010. CO-DEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

//www.hist-geo-space-
sci.net/1/45/2010/.

Schwach:2013:BRF

[Sch13] V. Schwach. Book review:

“The Fluid Envelope of our Planet: How the Study of Ocean Currents Became a Science”. History of Geoand Space Sciences, 4(2): 105–106, ???? 2013. CO-DEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

//www.hist-geo-space-
sci.net/4/105/2013/hgss-
4-105-2013.pdf.

Schmidtke:2015:EUS

[Sch15] G. Schmidtke. Extreme ultraviolet spectral irradiance measurements since 1946. History of Geoand Space Sciences, 6(1): 3–22, ???? 2015. CO-DEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

//www.hist-geo-space-
sci.net/6/3/2015/.

Schweitzer:2023:FDG

[Sch23] Johannes Schweitzer. Foundation of the Deutsche Geophysikalische Gesellschaft in its international context. History of Geo- and Space Sciences, 14(1):15– 22, ???? 2023. CO-DEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

//hgss.copernicus.org/
articles/14/15/2023/.

Schlegel:2024:LTE

[Sch24] Kristian Schlegel. Lightning and thunder explanations in encyclopedias — from ancient Greece to Wikipedia. History of Geoand Space Sciences, 15(2): 71–80, ???? 2024. CO-DEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

//hgss.copernicus.org/
articles/15/71/2024/.

Schlegel:2011:HGA

[SG11] K. Schlegel and G. Gregori. “Historian of geomagnetism and aeronomy”: Obituary — Dr. Wilfried Schröder. History of Geoand Space Sciences, 2(1):

83–84, ???? 2011. CO-DEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

//www.hist-geo-space-
sci.net/2/83/2011/.

Shcheglov:2018:LCP

[Shc18] Dmitry A. Shcheglov.

The length of coastlines in Ptolemy’s Geography and in ancient periploi. History of Geoand Space Sciences, 9(1): 9–24, ???? 2018. CO-DEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

//www.hist-geo-space-
sci.net/9/9/2018/.

Shcheglov:2020:CPE

[Shc20] Dmitry A. Shcheglov. The configuration of the Pontus Euxinus in Ptolemy’s geography. History of Geoand Space Sciences, 11(1): 31–51, ???? 2020. CO-DEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

//www.hist-geo-space-
sci.net/11/31/2020/.

Sheeley:2017:OWS

[She17] Neil R. Sheeley, Jr. Origin of the Wang–Sheeley–Arge solar wind model. History of Geo- and Space Sciences, 8(1):21–28, ???? 2017. CODEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL

https://www.hist-geo-
space-sci.net/8/21/2017/

.

Silverman:2012:BRH

[Sil12] S. Silverman. Book review: “Harald Moltke — Painter of the Aurora”. History of Geo- and Space Sciences, 3(1):127–129, ???? 2012. CODEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

//www.hist-geo-space-
sci.net/3/127/2012/hgss-
3-127-2012.pdf.

Schlegel:2014:WSE

[SL14] K. Schlegel and H. Lühr. Willy Stoffregen — an early pioneer of advanced ionospheric and auroral research. History of Geoand Space Sciences, 5(2): 149–154, ???? 2014. CO-DEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

//www.hist-geo-space-
sci.net/5/149/2014/.

Schweitzer:2019:IOP

[SL19] Johannes Schweitzer and Thorne Lay. IASPEI: its origins and the promotion of global seismology. History of Geo- and Space Sciences, 10(1):173–180, ???? 2019. CODEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL

https://www.hist-geo-
space-sci.net/10/173/
2019/.

Smallwood:2018:APV

[Sma18] John R. Smallwood. The attraction of the pyramids: virtual realization of Hutton’s suggestion to improve Maskelyne’s 1774 Earth density estimate. History of Geo- and Space Sciences, 9(1):1–7, ???? 2018. CODEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

//www.hist-geo-space-
sci.net/9/1/2018/.

Sano:2021:EHS

[SN21] Yasuharu Sano and Hiroshi Nagano. Early history of sudden commencement investigation and some newly discovered historical facts. History of Geo- and Space Sciences, 12(2):131–162, ???? 2021. CODEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

//hgss.copernicus.org/
articles/12/131/2021/.

Soffel:2015:HMM

[Sof15]  H. C. Soffel.  History

of the Munich–Maisach– Fürstenfeldbruck Geomagnetic Observatory. History of Geo- and Space Sciences, 6(2):65–86, ???? 2015. CODEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL

https://www.hist-geo-
space-sci.net/6/65/2015/

.

Sperberg:2021:EEP

[Spe21] Ulrich Sperberg. Eduard Heis, an early pioneer in meteor research. History of Geo- and Space Sciences, 12(2):163–170, ???? 2021. CODEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

                                        //hgss.copernicus.org/
                                        articles/12/163/2021/.

Schlegel:2011:JCH

[SS11] K. Schlegel and S. Silverman. Johann Christian Heuson, a little-known auroral scholar of the early 18th century. History of Geo- and Space Sciences, 2 (2):89–95, ???? 2011. CO-DEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

                                        //www.hist-geo-space-
                                        sci.net/2/89/2011/.

Stauning:2011:DAS

[Sta11a] P. Stauning. Danish auroral science history. History of Geo- and Space Sciences, 2(1):1–28, ???? 2011. CODEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL

                                        https://www.hist-geo-
                                        space-sci.net/2/1/2011/

. See comment [Jør11] and reply [Sta11b].

Stauning:2011:RCD

[Sta11b]  P. Stauning.  Reply

to Comment on “Danish auroral science history” by P. Stauning in Hist. Geo Space Sci., 2, 1–28, 2011. History of Geoand Space Sciences, 2(2): 87, ???? 2011. CO-DEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

//www.hist-geo-space-
sci.net/2/87/2011/. See

[Sta11a, Jør11].

Stauning:2011:HMN

[Sta11c] Peter Stauning. Harald Moltke: nordlysets maler: Harald Moltkes malerier på Danmarks Meteorologiske Institut (Danish) Harald Moltke: painter of the aurora: Harald Moltke’s paintings at the Danish Meteorological Institute. Forlaget Epsilon.dk, Frederiksberg, Danmark, 2011. ISBN 87-993384-3-2. 216 pp. LCCN ND723.M55 S83 2011.

Sudiro:2014:EET

[Sud14] P. Sudiro. The Earth expansion theory and its transition from scientific hypothesis to pseudoscientific belief. History of Geoand Space Sciences, 5(1): 135–148, ???? 2014. CO-DEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

//www.hist-geo-space-
sci.net/5/135/2014/. See

comment [Edw16].

Smythe-Wright:2019:ITO

[SWGM^{+}19] Denise Smythe-Wright, W. John Gould, Trevor J. McDougall, Stefania Sparnocchia, and Philip L. Woodworth. IAPSO: tales from the ocean frontier. History of Geo- and Space Sciences, 10(1):137–150, ???? 2019. CODEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL

                                        https://www.hist-geo-
                                        space-sci.net/10/137/
                                        2019/.

Schroder:2010:GNG

[SWS10] W. Schröder, K.-H. Wiederkehr, and K. Schlegel. Georg von Neumayer and geomagnetic research. History of Geo- and Space Sciences, 1 (2):77–87, ???? 2010. CO-DEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

                                        //www.hist-geo-space-
                                        sci.net/1/77/2010/.

Taylor:2010:BRN

[Tay10] K. L. Taylor. Book review: “A Natural History of Time”. History of Geoand Space Sciences, 1(1): 43–44, ???? 2010. CO-DEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

                                        //www.hist-geo-space-
                                        sci.net/1/43/2010/hgss-
                                        1-43-2010.pdf.

Thrane:2018:HAR

[Thr18] Eivind V. Thrane. The history of Andøya Rocket Range. History of Geoand Space Sciences, 9(2): 141–156, ???? 2018. CO-DEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

//www.hist-geo-space-
sci.net/9/141/2018/.

Tordesillas:2024:HEG

[TPCNA24] Jose Manuel Tordesillas, Francisco Javier Pavón-Carrasco, Alberto Nuñez, and Ana Belén Anquela. Historical evolution of the geomagnetic declination at the Royal Observatory of Madrid. History of Geoand Space Sciences, 15(2): 81–94, ???? 2024. CO-DEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

//hgss.copernicus.org/
articles/15/81/2024/.

Valiuskevicius:2017:SKC

[Val17] Gintaras Valiuskevicius.

Steponas Kolupaila’s contribution to hydrological science development. History of Geo- and Space Sciences, 8(2):57–67, ???? 2017. CODEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL

https://www.hist-geo-
space-sci.net/8/57/2017/

.

Vaquero:2017:BLR

[Vaq17] José M. Vaquero. Ball lightning: a Renaissance account from Zafra (Spain). History of Geoand Space Sciences, 8(1): 53–56, ???? 2017. CO-DEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

                                        //www.hist-geo-space-
                                        sci.net/8/53/2017/.

Wannberg:2022:HEP

[Wan22] Gudmund Wannberg. History of EISCAT — part 5: Operation and development of the system during the first 2 decades. History of Geoand Space Sciences, 13(1): 1–21, ???? 2022. CO-DEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

                                        //hgss.copernicus.org/
                                        articles/13/1/2022/.

Weiss:2016:HJI

[Wei16] J. Weiß. History of the Juliusruh ionospheric observatory on Rügen. History of Geo- and Space Sciences, 7(1):1–22, ???? 2016. CODEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL

                                        https://www.hist-geo-
                                        space-sci.net/7/1/2016/

.

Woodman:2019:EHJ

[WFBM19] Ronald F. Woodman, Donald T. Farley, Ben B. Balsley, and Marco A. Milla. The early history of the Jicamarca Radio Observatory and the incoherent scatter technique. History of Geo- and Space Sciences, 10(2):245–266, ???? 2019. CODEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL

https://www.hist-geo-
space-sci.net/10/245/
2019/.

Wilkinson:2015:BRC

[Wil15] P. Wilkinson. Book review:

Cosmogenic Radionuclides: Theory and Applications in the Terrestrial and Space Environments. History of Geo- and Space Sciences, 6 (1):1–2, ???? 2015. CO-DEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

//www.hist-geo-space-
sci.net/6/1/2015/.

Wilkinson:2018:PHR

[Wil18] Phil Wilkinson. Preface:

History of regional warning centers. History of Geo- and Space Sciences, 9(1):37, ???? 2018. CO-DEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

//www.hist-geo-space-
sci.net/9/37/2018/.

Winkler:2023:EMN

[Win23] Peter Winkler. The early meteorological network of the Societas Meteorologica Palatina (1781–1792): foundation, organization, and reception. History of Geo- and Space Sciences, 14(2):93–120, ???? 2023. CODEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

                                        //hgss.copernicus.org/
                                        articles/14/93/2023/.

Wittmann:2020:CFG

[Wit20] Axel D. Wittmann. Carl Friedrich Gauss and the Gauss Society: a brief overview. History of Geoand Space Sciences, 11(2): 199–205, ???? 2020. CO-DEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

                                        //hgss.copernicus.org/
                                        articles/11/199/2020/.

Wilkinson:2018:DAS

[WKC18] Phil Wilkinson, John A. Kennewell, and David Cole. The development of the Australian Space Forecast Centre (ASFC). History of Geoand Space Sciences, 9(1): 53–63, ???? 2018. CO-DEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

                                        //www.hist-geo-space-
                                        sci.net/9/53/2018/.

Woodworth:2020:TPM

[Woo20] Philip L. Woodworth. Tide prediction machines at the Liverpool Tidal Institute. History of Geoand Space Sciences, 11(1): 15–29, ???? 2020. CO-DEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

//www.hist-geo-space-
sci.net/11/15/2020/.

Woodworth:2018:TMJ

[WR18] Philip L. Woodworth and Glen H. Rowe. The tidal measurements of James Cook during the voyage of the Endeavour. History of Geo- and Space Sciences, 9 (1):85–103, ???? 2018. CO-DEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

//www.hist-geo-space-
sci.net/9/85/2018/.

Zolesi:2018:RCA

[ZC18] Bruno Zolesi and Ljiljana R. Cander. The role of COST Actions in unifying the European ionospheric community in the transition between the two millennia. History of Geoand Space Sciences, 9(1): 65–77, ???? 2018. CO-DEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https:

//www.hist-geo-space-
sci.net/9/65/2018/.

Zhang:2021:DGE

[ZW21] Zhihui Zhang and Rui Wang. The development of geophysics in the early period of the People’s Republic of China based on the Institute of Geophysics, Chinese Academy of Sciences (1950-1966). History of Geo- and Space Sciences, 12(1):21–41, ???? 2021. CODEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL

                                        https://hgss.copernicus.
                                        org/articles/12/21/2021/

.